OpenStudy (anonymous):

What percent of 90 is 18?

hero (hero):

\[\frac{is}{of}=\frac{\%}{100}\]

hero (hero):

\[\frac{18}{90}=\frac{\%}{100}\]

hero (hero):

Percent is unknown, so replace percent with x: \[\frac{18}{90}=\frac{x}{100}\] Cross multiply, solve for x

hero (hero):

90 x = 1800 x = 1800/90 x = 20 Therefore 20%

OpenStudy (anonymous):

ok thx so much it really helped :)

hero (hero):

You can also re-write the beginning of the problem and separate components like so: what always gets replaced with x: x percent of 90 is 18

hero (hero):

and then x% of 90 is 18

OpenStudy (anonymous):

15% of what number is 24?

hero (hero):

But always remember this formula: \[\frac{is}{of}=\frac{\%}{100}\]

hero (hero):

For this one: 15% of x is 24 so: \[\frac{24}{x}=\frac{15}{100}\]

hero (hero):

Do you get it?

OpenStudy (anonymous):

not really wat goes in the x spot ?

hero (hero):

x is what you have to find

OpenStudy (anonymous):

oh ok let me work it out and see

hero (hero):

x is an unknown variable.

hero (hero):

the "of what number" is the part of the question that is unknown

OpenStudy (anonymous):

is it 36 ?